### v1

# FINALLY: fast and universal speech enhancement with studio-like quality

Abstract
In this paper, we address the challenge of speech enhancement in real-world recordings, which often contain various forms of distortion, such as background noise, reverberation, and microphone artefacts. We revisit the use of Generative Adversarial Networks (GANs) for speech enhancement and theoretically show that GANs are naturally inclined to seek the point of maximum density within the conditional clean speech distribution, which, as we argue, is essential for the speech enhancement task. We study various feature extractors for perceptual loss to facilitate the stability of adversarial training, developing a methodology for probing the structure of the feature space. This leads us to integrate WavLM-based perceptual loss into MS-STFT adversarial training pipeline, creating an effective and stable training procedure for the speech enhancement model. The resulting speech enhancement model, which we refer to as FINALLY, builds upon the HiFi++ architecture, augmented with a WavLM encoder and a novel training pipeline. Empirical results on various datasets confirm our model’s ability to produce clear, high-quality speech at 48 kHz, achieving state-of-the-art performance in the field of speech enhancement.
1 Introduction
Speech recordings are often contaminated with background noise, reverberation, reduced frequency bandwidth, and other distortions. Unlike classical speech enhancement (Ephraim & Malah, 1984; Pascual et al., 2017), which considers each task separately, universal speech enhancement (Serrà et al., 2022; Su et al., 2021; Liu et al., 2022) aims to restore speech from all types of distortions simultaneously. Thus, universal speech enhancement seeks to generalize across a wide range of distortions, making it more suitable for real-world applications where multiple distortions may coexist.
Recent studies have categorized the problem of speech enhancement as a task of learning the clean speech distribution conditioned on degraded signals (Lemercier et al., 2023; Serrà et al., 2022; Richter et al., 2023). This problem is often addressed using diffusion models (Ho et al., 2020; Song et al., 2020), which are renowned for their exceptional ability to learn distributions. Diffusion models have recently achieved state-of-the-art results in universal speech enhancement (Serrà et al., 2022). However, the impressive performance of diffusion models comes with the high computational cost of their iterative inference process.
It is important to note that the speech enhancement problem does not require the model to learn the entire conditional distribution. In practice, when presented with a noisy speech sample, the goal is often to obtain the most probable clean speech sample that retains the lexical content and voice of the original. This contrasts with applications such as text-to-image synthesis (Ramesh et al., 2022; Lee et al., 2024; Rombach et al., 2022), where the objective is to generate a variety of images for each text prompt due to the higher level of uncertainty and the need for diverse options to select the best image. For most speech enhancement applications, such as voice calls and compensation for poor recording conditions, capturing the entire conditional distribution is not necessary. Instead, it is more important to retrieve the most likely sample of this distribution (the main mode), which might be a simpler task.
Diffusion models’ main advantage over generative adversarial networks (GANs) (Goodfellow et al., 2014) is their ability to capture different modes of the distribution. However, we argue that this property is not typically required for the task of speech enhancement and may unnecessarily complicate the operation of the neural network. Conversely, we show that GANs tend to retrieve the main mode of the distribution—precisely what speech enhancement should typically do.
Therefore, in this work, we revisit the GAN framework for speech enhancement and demonstrate that it provides rapid and high-quality universal speech enhancement. Our model outperforms both diffusion models and previous GAN-based models, achieving an unprecedented level of quality on both simulated and real-world data.
Our main contributions are as follows:
  1. 1.
    We theoretically analyse the adversarial training with the least squares GAN (LS-GAN) loss and demonstrate that a generator predicting a single sample per input condition (producing a conditional distribution that is a delta function) is incentivized to select the point of maximum density. Therefore, we establish that LS-GAN training can implicitly regress for the main mode of the distribution, aligning with the objectives of the speech enhancement problem.
  2. 2.
    We investigate various feature extractors as backbones for perceptual loss and propose criteria for selecting an extractor based on the structure of its feature space. These criteria are validated by empirical results from a neural vocoding task, indicating that the convolutional features of the WavLM neural network(Chen et al., 2022b) are well-suited for perceptual loss in speech generation.
  3. 3.
    We develop a novel model for universal speech enhancement that integrates the proposed perceptual loss with MS-STFT discriminator training (Défossez et al., 2023) and enhances the architecture of the HiFi++ generator (Andreev et al., 2022) by combining it with a self-supervised pre-trained WavLM encoder (Chen et al., 2022b). Our final model delivers state-of-the-art performance on real-world data, producing high-quality, studio-like speech at 48 kHz.
2 Mode Collapse and Speech Enhancement
The first question that we address is what is the practical purpose of a speech enhancement model. The practical goal of a speech enhancement model is to restore the audio signal containing the speech characteristics of the original recording, including the voice, linguistic content, and prosody. Thus, loosely speaking, the purpose of the speech enhancement task for many applications is not “generative” in its essence, in the sense that the speech enhancement model should not generate new speech content but rather “refine” existing speech as if it was recorded in ideal conditions (studio-like quality). From the mathematical point of view, this means that the speech enhancement model should retrieve the most probable reconstruction of the clean speech $y$ given the corrupted version $x$, i.e., $y=\mathrm{arg\,max}_{y}\ p_{\text{clean}}(y|x)$.
This formulation re-considers the probabilistic speech enhancement formulation, which is widely used in the literature. In such formulation, the speech enhancement model is aimed to capture the entire conditional distribution $p_{\text{clean}}(y|x)$. This formulation might be especially appealing in situations with high generation ambiguity, e.g., a low SNR scenario where clean speech content could not be restored unambiguously. In this case, the speech enhancement model could be used to generate multiple reconstructions, the best of which is then selected by the end user. However, we note that this formulation might be redundant and not particularly relevant for many practical applications since the ambiguity in generation can be resolved by more straightforward means such as conditioning on linguistic content (Koizumi et al., 2023c).
In practice, for many applications, a more natural way of formalizing speech enhancement is to treat it as a regression problem which aims at predicting the point of highest probability of the conditional distribution $\mathrm{arg\,max}_{y}\ p_{\text{clean}}(y|x)$. This formulation has the advantage of simplifying the task, since finding the highest mode of the distribution might be significantly easier than learning the entire distribution. Therefore, the speech enhancement models built for this formulation are likely to be more efficient after deployment since they solve a simpler task. We note that in the context of speech enhancement, the speed of inference is always of major concern in practice.
Given this formulation, we argue that the framework of generative adversarial networks (GANs) is more naturally suited for the speech enhancement problem than diffusion models. We show that GAN training naturally leads to the mode-seeking behaviour of the generator, which aligns with the formulation introduced above. Additionally, GANs enjoy one forward pass inference, which is in contrast to the iterative nature of diffusion models.
Let $p_{g}(y|x)$ be a family of waveform distributions produced by the generator $g_{\theta}(x)$. Mao et al. (2017) showed that training with Least Squares GAN (LS-GAN) leads to the minimization of the Pearson $\chi^{2}$ divergence $\chi^{2}_{\text{Pearson}}\left(p_{g}\middle\|\frac{p_{\text{clean}}+p_{g}}{2}\right)$. We propose that if $p_{g}(y|x)$ approaches $\delta(y-g_{\theta}(x))$ under some parametrization, the minimization of this divergence leads to $g_{\theta}(x)=\mathrm{arg\,max}_{y}\ p_{\text{clean}}(y|x)$. This means that if the generator deterministically predicts the clean waveform from the degraded signal, the LS-GAN loss encourages the generator to predict the point of maximum $p_{\text{clean}}(y|x)$ density. We note that although prior work by (Li & Farnia, 2023) demonstrated the mode-covering property for the optimization of Pearson $\chi^{2}$ divergence, our result pertains to a deterministic generator setting, which is outside the scope of analysis provided by Li & Farnia (2023).
[…]
Proposition 1.
Let $p_{\text{clean}}(y|x)>0$ be a finite and Lipschitz continuous density function with a unique global maximum and $p^{\xi}_{g}(y|x)=\xi^{n}/2^{n}\cdot\mathbf{1}_{y-g_{\theta}(x)\in[-1/\xi,1/\xi]^{n}}$, then
[…]
Thus, LS-GAN training under ideal conditions should lead to the solution $g_{\theta}(x)=\mathrm{arg\,max}_{y}\ p_{\text{clean}}(y|x)$ for the generator. In practice, however, success is highly dependent on technicalities, such as additional losses to stabilize training and architectures of neural networks. Below, we address these questions by revisiting the notion of perceptual loss for audio generation and assessing the effectiveness of neural architectures.
3 Perceptual Loss for Speech Generation
Adversarial training is known for its instability issues (Brock et al., 2018). It often leads to suboptimal solutions, mode collapse, and gradient explosions. For paired tasks, including speech enhancement, adversarial losses are often accompanied by additional regressive losses to stabilize training and guide the generator towards useful solutions (Kong et al., 2020; Su et al., 2020). In the context of GAN mode-seeking behaviour discussed above, regressive losses could be seen as a merit to push the generator towards the “right” (most-probable) mode. Therefore, finding an appropriate regression loss to guide adversarial training is of significant importance.
Historically, initial attempts to apply deep learning methods to speech enhancement were based on treating this problem as a predictive task (Defossez et al., 2020; Hao et al., 2021; Chen et al., 2022a; Isik et al., 2020). Following the principle of empirical risk minimization, the goal of predictive modelling is to find a model with minimal average error over the training data. Given a noisy waveform or spectrogram, these approaches attempt to predict the clean signal by minimizing point-wise distance in waveform and spectrum domains or jointly in both domains, thus treating this problem as a predictive task. However, given the severe degradations applied to the signal, there is an inherent uncertainty in the restoration of the speech signal (i.e., given the degraded signal, the clean signal is not restored unambiguously), which often leads to oversmoothing (averaging) of the predicted speech. A similar phenomenon is widely known in computer vision (Ledig et al., 2017).
One promising idea to reduce the averaging effect is to choose an appropriate representation space for regression, which is less “entangled” than waveform or spectrogram space. In simpler terms, the regression space should be designed so that averaged representation of sounds that are indistinguishable to humans (such as the same phoneme spoken by the same speaker with the same prosody) is still representation of this sound (see Section A.1).
We formulate two heuristic rules to compare different regression spaces based on their structure:
  • Clustering rule: Representations of identical speech sounds should form one cluster that is separable from clusters formed by different sounds.
  • SNR rule: Representations of speech sounds contaminated by different levels of additive noise should move away from the cluster of clean sounds monotonically with the increase in the noise level.
The clustering rule ensures that minimizing the distance between samples in the feature space causes the samples to correspond to the same sound. The SNR rule ensures that minimizing the distance between features does not contaminate the signal with noise, meaning noisy samples are placed distantly from clean samples.
Figure 1: Illustration of heuristic rules for feature space structure. The Clustering rule (left) states that representations of the same speech sound should form clusters. The SNR rule (right) states that noise samples should deviate from the centre of the cluster as the amount of noise increases. Illustrations created using real samples are presented in  Figure 8,  Figure 7
In practice, we check these conditions by the following procedure:
  1. 1.
    We sample identical speech sounds with a multi-speaker VITS text-to-speech model (Kim et al., 2021). We define the group of waveforms corresponding to the same sound as 0.5-second waveform segments generated by the same speaker with the same text and the same phoneme duration (i.e., the stochastic duration predictor is used only once for each group). Thus, the waveform variation is produced due to sampling of latents from the prior distribution. We note that we do not explicitly fix the prosody; however, we observed that prosody variation was low by default and the sounds generated by this procedure were mostly perceptually indistinguishable from each other while corresponding to different waveform realizations. Overall, we define 354 groups of sounds with 80 speakers saying 177 different phrases, 2 different speakers for each phrase; we sample 20 samples for each group, totalling 7080 waveform segments corresponding to 354 groups.
  2. 2.
    Each waveform segment is then mapped to the feature space, and the resulting tensors are flattened to obtain a vector for each sample. The vectors are clustered using K-means clustering (MacQueen et al., 1967), the number of clusters is set to the number of groups (354). After clustering, we compute the Rand index (Rand, 1971) of the clustering produced by K-means with the clustering induced by initial groups of sounds splits. We treat the resulting number as a way to measure the adherence to the clustering rule.
  3. 3.
    Each waveform segment within a group is randomly mixed with noise at SNR levels ranging from 10 to 20 dB. The segments are then mapped to the feature space to obtain a vector for each sample. For each noisy sample, we compute the Euclidean distance between its features and the centre of the clean feature cluster. After that, we compute the negative Spearman’s rank correlation between the SNR level and the distance from the sample to the centre of the cluster. The negative correlations are averaged over all groups of sounds, and the resulting number is treated as a quantitative measure of adherence to the SNR rule.
Using these metrics, we assess the effectiveness of different feature spaces formed by several speech feature extractors, as well as conventional representations of speech. Namely, we produce features by Wav2Vec 2.0 (Baevski et al., 2020), WavLM (Chen et al., 2022b), the encoder of EnCodec (Défossez et al., 2023), and CDPAM (Manocha et al., 2021). As a conventional representation, we use waveform and spectrogram features. For Wav2Vec 2.0 and WavLM, we consider the output of the last transformer layer and the output of the convolutional encoder as separate cases. We also train a HiFi-GAN generator (Kong et al., 2020) with each feature type used as a representation for mean squared error loss computation on a neural vocoding task. To assess experimentally the suitability of the feature space to be used as a loss function for waveform generation, we report MOS scores for samples generated by vocoders trained with each feature map. The results are presented in Table 1.
Table 1: Comparison of different features using Clustering rule, SNR rule, and MOS on neural vocoding.
Feature
space
Rand score ($\uparrow$)
(Clustering rule)
Negative correlation ($\uparrow$)
(SNR rule)
MOS ($\uparrow$)
Waveform $00.00\pm 0$ 0.00 $0.31\pm 0.02$ Failed
Spectrogram $00.00\pm 0$ 0.00 $0.08\pm 0.03$ $1.78\pm 0.08$
Wav2Vec 2.0 $0.25\pm 0.03$ $0.19\pm 0.03$ $1.65\pm 0.08$
Wav2Vec 2.0-conv $0.94\pm 0.01$ $0.78\pm 0.02$ $2.23\pm 0.09$
WavLM $0.46\pm 0.05$ $0.46\pm 0.03$ $1.71\pm 0.07$
WavLM-conv $\mathbf{0.96\pm 0.01}$ $\mathbf{0.89\pm 0.02}$ $\mathbf{3.27\pm 0.10}$
EnCodec $0.55\pm 0.03$ $0.67\pm 0.03$ $1.80\pm 0.08$
CDPAM $00.00\pm 0$ 0.00 $0.17\pm 0.03$ Failed
As can be seen, the features extracted by the convolutional encoder of the WavLM model present the best performance according to both our internal metrics and MOS score on the neural vocoding task. We have also tested the other layers of WavLM but did not observe significant improvements when using other layers (see Section A.4). Empirically, we have found that if WavLM-conv feature loss is used together with spectrogram loss (L1-distance magnitudes of STFT), the performance on the vocoding task increases significantly, likely due to the looseness of the WavLM representation (Section A.1). Without any adversarial training, the WavLM-conv+L1-STFT loss (LMOS-loss, Equation 2) achieves an MOS score of $4.31\pm 0.08$, approaching the original adversarially trained HiFi GAN generator which achieves $4.69\pm 0.05$ (Section A.3).
4 FINALLY
4.1 Architecture
Our method is based on HiFi++ (Andreev et al., 2022). The HiFi++ generator is a four-component neural network consisting of SpectralUNet, Upsampler, WaveUNet, and SpectralMaskNet modules. SpectralUNet is responsible for initial preprocessing of audio in the spectral domain using two-dimensional convolutions. The Upsampler is a HiFi-GAN generator-based module that increases the temporal resolution of the input tensor, mapping it to the waveform domain. WaveUNet performs post-processing in the waveform domain and improves the output of the Upsampler by incorporating phase information gleaned directly from the raw input waveform. Finally, SpectralMaskNet is applied to perform spectrum-based post-processing and, thus, remove any possible artefacts that remained after WaveUNet. Thus, the model alternates between time and frequency domains, allowing for effective audio restoration.
We introduce two modifications to the HiFi++ generator’s architecture. First, we modify the generator by incorporating WavLM-large model output (last hidden state of the transformer) as an additional input to the Upsampler. Prior works (Hung et al., 2022; Byun et al., 2023) have demonstrated the usefulness of Self-Supervised Learning (SSL) features for speech enhancement tasks, and we validate this by observing significant performance gains from using SSL features. Second, we introduce the Upsample WaveUNet at the end of the generator. This module acts as a learnable upsampler of the signal sampling rate. For its architecture, we use the WaveUNet with an additional convolutional upsampling block in the decoder that upsamples the temporal resolution by 3 times. This allows the model to output a 48 kHz signal while taking a 16 kHz signal as input.
Figure 2: FINALLY model architecture.
4.2 Data and training
We use LibriTTS-R (Koizumi et al., 2023b), DAPS-clean (Mysore, 2014) as the sources of clean speech data. LibriTTS-R is used at 16 kHz, while DAPS at 48 kHz. Noise samples were taken from the DNS dataset (Dubey et al., 2022). After mixing with noise, we apply several digital distortion effects (see Footnote 2 for details).
We train the model in three stages. The first two stages concentrate on restoring the original speech content, and the final stage aims to enhance the aesthetic perception of the speech. The multi-stage approach is necessary due to the characteristics of the employed datasets: the LibriTTS-R dataset has a lot of samples but limited perceptual quality, whereas the DAPS dataset is high-quality but contains a smaller number of samples. Consequently, we utilize the LibriTTS-R dataset for learning speech content restoration and the DAPS dataset for aesthetic stylization.
The loss functions that we use can be written as follows:
$\mathcal{L}_{\text{LMOS}}(\theta)=\underset{x,y\sim p(x,y)}{\mathbb{E}}\left[100\cdot\|\phi(y)-\phi(g_{\theta}(x))\|_{2}^{2}+\||\mathrm{STFT}(y)|-|\mathrm{STFT}(g_{\theta}(x))|\|_{1}\right],$ (2)
$\mathcal{L}_{\text{gen}}(\theta)=\underbrace{\overbrace{\underbrace{\lambda_{\text{LMOS}}\cdot\mathcal{L}_{\text{LMOS}}(\theta)}_{\text{1st stage (16 kHz)}}+\lambda_{\text{GAN}}\cdot\mathcal{L}_{\text{GAN-gen}}(\theta)+\lambda_{\text{FM}}\cdot\mathcal{L}_{\text{FM}}(\theta)}^{\text{2ndstage (16 kHz)}}+\lambda_{\text{HF}}\cdot\mathcal{L}_{\text{HF}}(\theta)}_{\text{3rd stage (48 kHz)}},$ (3)
$\mathcal{L}_{\text{disc}}(\varphi_{i})=\mathcal{L}_{\text{GAN-disc}}(\varphi_{i}),\quad i=1,\ldots,k.$ (4)
Here, $\phi$ denotes the WavLM-Conv feature mapping, $g_{\theta}(x)$ denotes the generator neural network with parameters $\theta$, $\mathcal{L}_{\text{GAN-gen}}(\theta)$ denotes the LS-GAN generator loss (Mao et al., 2017), $\mathcal{L}(\theta)$ denotes the combined generator loss, $\mathcal{L}_{\text{GAN-disc}}(\varphi_{i})$ denotes the LS-GAN discriminator (Mao et al., 2017) loss for the $i$-th discriminator with parameters $\varphi_{i}$, $\mathcal{L}_{\text{FM}}$ denotes the feature matching loss (Kumar et al., 2019; Défossez et al., 2023), $\mathcal{L}_{\text{HF}}$ denotes the human feedback loss, and $*\lambda_{\textbf{*}}$ denotes the corresponding loss weights. Following Défossez et al. (2023), we employ 5 discriminators with STFT window lengths of $[2048,1024,512,256,128]$ for the 2nd stage and 5 discriminators with STFT window lengths of $[4096,2048,1024,512,256]$ for the 3rd stage, which allows for the capture of spectral information at different resolutions.
Stage 1: Firstly, we train the FINALLY architecture without Upsample WaveUNet (we refer to this truncated architecture as FINALLY-16) at a 16 kHz sampling rate on the LibriTTS-R dataset. We train the model with the proposed LMOS regression loss to provide the generator with a better initialization before adversarial training.
Stage 2: Second, we start adversarial training of FINALLY-16 with MS-STFT discriminators (Défossez et al., 2023) and the LMOS loss. The learning of the generator undergoes a linear warm-up to give the discriminators time to learn meaningful representations. At this stage, we focus the generator on producing a reliable reconstruction of linguistic content by assigning larger values for LMOS and feature matching losses ($\lambda_{\text{GAN-gen}}=0.4$, $\lambda_{\text{FM}}=20$, $\lambda_{\text{LMOS}}=20$).
Stage 3: Lastly, we attach Upsample WaveUNet to FINALLY-16 and start adversarial training of the FINALLY model to produce 48 kHz output. We subsample 48 kHz waveforms of the DAPS dataset and apply distortions to form the 16 kHz input. The discriminators are initialized randomly, and the learning rate for the generator undergoes a warm-up similar to the second stage. This stage is focused on producing the final output; therefore, we focus the generator on perceptual quality by increasing the relative weight of GAN loss ($\lambda_{\text{GAN-gen}}=5$, $\lambda_{\text{FM}}=15$, $\lambda_{\text{LMOS}}=0.5$) and introducing additional human feedback loss $\mathcal{L}_{\text{HF}}$, which is based on UTMOS and PESQ metrics. The UTMOS loss (Saeki et al., 2022) is based on a neural network model that simulates subjective MOS metric results. On the other hand, the PESQ loss [^1] https://github.com/audiolabs/torch-pesq delivers a differentiable version of the PESQ metric (Rix et al., 2001b), as presented by Kim et al. (2019); Martin-Donas et al. (2018). Both UTMOS and PESQ help enhance speech aesthetic quality, as they incorporate insights from human preference studies. These metrics are differentiable with respect to their inputs, making them suitable for use as loss functions (multiplied by negative constants) $\mathcal{L}_{\text{HF}}=-20\cdot\mathcal{L}_{\text{UTMOS}}-2\cdot\mathcal{L}_{\text{PESQ}}$, $\lambda_{\text{HF}}=1$.
5 Related Work
Self-supervised features for speech enhancement
Several works have employed self-supervised (SSL) features  (Baevski et al., 2020; Chen et al., 2022b; Hsu et al., 2021) for training of speech enhancement models as intermediate representation (Wang et al., 2024; Koizumi et al., 2023c), auxiliary space for loss computation (Sato et al., 2023; Hsieh et al., 2020; Close et al., 2023a, b) or input features (Hung et al., 2022; Byun et al., 2023). Sato et al. (2023); Close et al. (2023a, b); Hsieh et al. (2020) proposed to use features of self-supervised models as an auxiliary space for regression loss computation. In our work, we similarly study the effect of SSL features on speech enhancement models; however, our study systematically compares different feature backbones and develops criteria for probing the structure of different feature spaces. In particular, we show that features of the WavLM’s convolutional encoder are the most effective for the loss function, while recent work (Sato et al., 2023) has used the outputs of transformer layers. Close et al. (2023a, b) proposed a similar loss function for speech enhancement based on convolutional features of HuBERT (Hsu et al., 2021), our work can be considered as extension of this work as we show that additional spectrogram loss greatly benefits the quality and provide experimental evidence for the advantages of LMOS-loss usage for adversarial training.
GAN-based approaches
The HiFi GAN works (Su et al., 2020, 2021) consider a GAN-based approach to speech enhancement, which is similar to ours. Importantly, these works base their generator architectures on feed-forward WaveNet (Rethage et al., 2018), a fully-convolutional neural network that operates at full input resolution, leading to slow training and inference times. We show that our model is able to achieve superior quality compared to this model while being much more efficient.
Koizumi et al. (2023c) proposes to use w2v-BERT features (Chung et al., 2021) as intermediate representations for speech restoration. The features extracted from the noisy waveform are processed by a feature cleanser which is conditioned on the text representation extracted from transcripts via PnG-BERT, and speaker embedding extracted from the waveform. The feature cleanser is trained to minimize a regression loss between w2v-BERT features extracted from the clean signal and the predicted ones. At the second stage, the WaveFit vocoder is trained to synthesize a waveform based on the predicted features (Koizumi et al., 2023a). An important difference with our work is that the method uses a text transcript, while our model does not. Despite this difference, we show that our model delivers better perceptual quality as reported by human listeners (Appendix B).
Diffusion-based approaches
The recent success of diffusion-based generative models has led to their use in a wide array of applications, including speech enhancement. Numerous studies (Lemercier et al., 2023; Welker et al., 2022; Richter et al., 2023; Lay et al., 2023) have applied diffusion models in various configurations to generate high-fidelity, clear speech from noisy input. For instance, Welker et al. (2022); Richter et al. (2023) introduced a novel stochastic diffusion process to design a generative model for speech enhancement in the complex STFT domain for speech denoising and dereverberation. In the UNIVERSE study (Serrà et al., 2022), the authors propose a diffusion model for universal speech enhancement. They create a paired dataset using 55 different distortions and train a conditional diffusion model on it. Although their model performs well in terms of quality, it requires up to 50 diffusion steps to produce the final audio. The authors demonstrate that the number of steps can be reduced; however, the effect of this reduction on perceptual quality remains somewhat unclear. In our experiments, we show that our model can achieve similar results with just a single forward pass.
6 Results
6.1 Evaluation
We evaluate our models using the following sources of data:
VoxCeleb Data: 50 audio clips selected from VoxCeleb1 (Nagrani et al., 2017) to cover the Speech Transmission Index (STI) range of 0.75-0.99 uniformly and balanced across male and female speakers.
UNIVERSE Data: 100 audio clips randomly generated by the UNIVERSE (Serrà et al., 2022) authors from clean utterances sampled from VCTK and Harvard sentences, together with noises/backgrounds from DEMAND and FSDnoisy18k. The data contains various artificially simulated distortions including band limiting, reverberation, codec, and transmission artefacts. Please refer to (Serrà et al., 2022) for further details.
The validation data of UNIVERSE was artificially simulated from clean speech recordings using the same pipeline that the authors utilized for training. Therefore, we must note that the comparison is conducted in a manner advantageous to UNIVERSE, since our data simulation pipeline is different.
VCTK-DEMAND: we use validation samples from popular Valentini denoising benchmark (Valentini-Botinhao et al., 2017). This dataset is used for a broad comparison with a wide range of speech enhancement models. The test set (824 utterances) includes artificially simulated noisy samples from 2 speakers with 4 SNR (17.5, 12.5, 7.5, and 2.5 dB).
We utilize DNSMOS (Reddy et al., 2022), UTMOS (Saeki et al., 2022), and WV-MOS (Andreev et al., 2022) as non-intrusive metrics to objectively assess the samples generated by our speech enhancement model on the VoxCeleb dataset. The non-intrusive nature of these metrics is essential since the dataset comprises recordings from real-life scenarios, lacking ground-truth samples.
In addition, we compute the Phoneme Error Rate (PhER) and Word Error Rate (WER) by comparing ground truths with the generated samples (see Appendix D for details) for both the UNIVERSE and VCTK-DEMAND datasets. For subjective quality assessment, we conduct 5-point Mean Opinion Score (MOS) tests. All audio clips are normalized to ensure volume differences do not influence the raters’ evaluations. The raters are required to be English speakers using appropriate listening equipment (more details in Figure 10).
The Real-Time Factor (RTF) is determined by measuring the processing time (in seconds) for a 10-second audio segment on a V100 GPU and then dividing this time by 10. All confidence intervals are calculated using bootstrapping method.
Further, we evaluate our model on the VCTK-DEMAND dataset (Valentini-Botinhao et al., 2017) with additional metrics such as PESQ (Rix et al., 2001a), STOI (Taal et al., 2011), and SI-SDR (Roux et al., 2018). These metrics are included to ensure consistent comparison with previous works.
6.2 Comparison with existing approaches
Table 2: Comparison with prior work on Voxceleb and UNIVERSE validation data.
VoxCeleb (HiFi-GAN-2 validation set, real data)
Model MOS ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) DNSMOS ($\uparrow$) - RTF ($\downarrow$)
Input 3.46 $\pm$ 0.07 2.76 $\pm$ 0.13 2.90 $\pm$ 0.16 2.72 $\pm$ 0.11 - -
VoiceFixer 3.41 $\pm$ 0.07 2.60 $\pm$ 0.09 2.79 $\pm$ 0.09 3.08 $\pm$ 0.06 - 0.02
DEMUCS 3.79 $\pm$ 0.07 3.51 $\pm$ 0.08 3.72 $\pm$ 0.08 3.27 $\pm$ 0.04 - 0.08
STORM 3.75 $\pm$ 0.06 3.29 $\pm$ 0.08 3.54 $\pm$ 0.09 3.17 $\pm$ 0.04 - 1.05
BBED 3.97 $\pm$ 0.06 3.30 $\pm$ 0.10 3.47 $\pm$ 0.08 3.23 $\pm$ 0.04 - 0.43
HiFi-GAN-2 4.47 $\pm$ 0.05 3.67 $\pm$ 0.09 3.96 $\boldsymbol{\pm}$ 0.06 3.32 $\boldsymbol{\pm}$ 0.03 - 0.50
Ours 4.63 $\boldsymbol{\pm}$ 0.04 4.05 $\boldsymbol{\pm}$ 0.07 3.98 $\boldsymbol{\pm}$ 0.06 3.31 $\boldsymbol{\pm}$ 0.04 - 0.03
UNIVERSE validation set (simulated data)
Model MOS ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) DNSMOS ($\uparrow$) PhER ($\downarrow$) RTF ($\downarrow$)
Input 2.87 $\pm$ 0.05 2.27 $\pm$ 0.28 1.72 $\pm$ 0.61 2.25 $\pm$ 0.19 0.31 $\pm$ 0.05 -
Ground Truth 4.39 $\pm$ 0.05 4.26 $\pm$ 0.06 4.28 $\pm$ 0.06 3.33 $\pm$ 0.04 0 -
UNIVERSE 4.10 $\pm$ 0.07 3.89 $\pm$ 0.15 3.85 $\pm$ 0.12 3.23 $\boldsymbol{\pm}$ 0.07 0.20 $\pm$ 0.04 0.5
Ours (16 kHz) 3.99 $\pm$ 0.07 4.21 $\boldsymbol{\pm}$ 0.10 4.43 $\boldsymbol{\pm}$ 0.07 3.25 $\boldsymbol{\pm}$ 0.05 0.14 $\boldsymbol{\pm}$ 0.03 0.03
Ours 4.23 $\boldsymbol{\pm}$ 0.07 4.21 $\boldsymbol{\pm}$ 0.10 4.43 $\boldsymbol{\pm}$ 0.08 3.25 $\boldsymbol{\pm}$ 0.05 0.14 $\boldsymbol{\pm}$ 0.03 0.03
Table 3: Comparison with the baselines on VCTK-DEMAND.
Model MOS ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) DNSMOS ($\uparrow$) PESQ ($\uparrow$) STOI ($\uparrow$) SI-SDR ($\uparrow$) WER ($\downarrow$)
Input $3.18\pm 0.07$ $2.62\pm 0.16$ $2.99\pm 0.24$ $2.53\pm 0.10$ $1.98\pm 0.17$ $0.92\pm 0.01$ $8.4\pm 1.2$ $0.09\pm 0.03$
MetricGAN+ $3.75\pm 0.06$ $3.62\pm 0.09$ $3.89\pm 0.10$ $2.95\pm 0.05$ $3.14\pm 0.10$ $0.93\pm 0.01$ $8.6\pm 0.7$ $0.10\pm 0.04$
DEMUCS $3.95\pm 0.06$ $3.95\pm 0.05$ $4.37\pm 0.06$ $3.14\pm 0.04$ $3.04\pm 0.12$ 0.95 $\boldsymbol{\pm}$ 0.01 $18.5\pm 0.6$ 0.07 $\boldsymbol{\pm}$ 0.03
HiFi++ $4.08\pm 0.05$ $3.89\pm 0.06$ $4.36\pm 0.06$ $3.10\pm 0.04$ $2.90\pm 0.12$ 0.95 $\boldsymbol{\pm}$ 0.01 $17.9\pm 0.6$ $0.08\pm 0.03$
HiFi-GAN-2 $4.13\pm 0.05$ $3.99\pm 0.05$ $4.26\pm 0.05$ $3.12\pm 0.05$ $3.14\pm 0.12$ 0.95 $\boldsymbol{\pm}$ 0.01 $18.6\pm 0.6$ 0.07 $\boldsymbol{\pm}$ 0.03
DB-AIAT $4.22\pm 0.05$ $4.02\pm 0.05$ $4.38\pm 0.06$ $3.18\pm 0.04$ 3.26 $\boldsymbol{\pm}$ 0.12 0.96 $\boldsymbol{\pm}$ 0.01 19.3 $\boldsymbol{\pm}$ 0.8 0.07 $\boldsymbol{\pm}$ 0.03
Ours (16 kHz) 4.41 $\boldsymbol{\pm}$ 0.04 4.32 $\boldsymbol{\pm}$ 0.02 4.87 $\boldsymbol{\pm}$ 0.05 3.22 $\boldsymbol{\pm}$ 0.04 $2.94\pm 0.10$ $0.92\pm 0.01$ $4.6\pm 0.3$ 0.07 $\boldsymbol{\pm}$ 0.03
Ours (48 kHz) 4.66 $\boldsymbol{\pm}$ 0.04 4.32 $\boldsymbol{\pm}$ 0.02 4.87 $\boldsymbol{\pm}$ 0.05 3.22 $\boldsymbol{\pm}$ 0.04 $2.94\pm 0.10$ $0.92\pm 0.01$ $4.6\pm 0.3$ 0.07 $\boldsymbol{\pm}$ 0.03
GT (16 kHz) $4.26\pm 0.05$ $4.07\pm 0.04$ $4.52\pm 0.04$ $3.16\pm 0.04$
GT (48 kHz) $4.56\pm 0.03$ $4.07\pm 0.04$ $4.52\pm 0.04$ $3.16\pm 0.04$
We consider BBED (Lay et al., 2023), STORM (Lemercier et al., 2023), and UNIVERSE (Serrà et al., 2022) diffusion models, along with Voicefixer and DEMUCS regression models, as our baselines. In addition, we consider our closest competitor, HiFi-GAN-2, as a GAN-based baseline. The data for comparison with HiFi-GAN-2 and UNIVERSE were taken from their demo pages, since the authors did not release any code. We conduct comparisons with BBED, STORM, Voicefixer, DEMUCS, and HiFi-GAN-2 on real-world VoxCeleb1 samples, and the comparison with UNIVERSE on the simulated data provided by the authors of this work. The results are presented in Table 2. We also compare these models in VCTK-DEMAND dataset, provided in Table 3. We complement this table by two additional models: MetricGAN+ (Fu et al., 2021) and DB-AIAT (Yu et al., 2021).
Importantly, our study confirms the observation from Serrà et al. (2022) that speech enhancers based on generative models significantly outperform regression-based approaches. Our model performs comparably in terms of perceptual MOS quality to all the considered baselines, while being more than five times as efficient as the closest competitors, HiFi-GAN-2 and UNIVERSE. We have also found that our model is less prone to hallucinating linguistic content than UNIVERSE, delivering a lower Phoneme Error Rate (PhER) value.
Lastly, we would like to comment on the results in Table 3. Our model outperforms baselines in subjective evaluation and no-reference metrics, e.g. UTMOS, but significantly in terms of PESQ (Rix et al., 2001a), STOI (Taal et al., 2011) and SI-SDR (Roux et al., 2018). However, there have been numerous works consistently reporting low correlation of reference-based metrics with human perceptual judgment (Manocha et al., 2022; Manjunath, 2009; Andreev et al., 2022). In particular, the study (Manocha et al., 2022) reports that no-reference metrics (including DNSMOS, reported in our work) correlate significantly better with human perception and therefore have higher relevance for objective comparison between methods. Furthermore, in our study, we report the MOS score, which directly reflects human judgments of restoration quality.
6.3 Ablation study
To validate the improvements proposed in this work, we conduct an ablation study assessing the effectiveness of the design choices made.
Firstly, we compare the LMOS loss against two other regression losses in the context of training a small speech enhancement model (10 times smaller than the final model; please see the Section D.2 for details). The first regression loss is the Mel-Spectrogram loss, which was proposed by Kong et al. (2020). As the second alternative, we employ the Reconstruction loss (RecLoss) proposed by Défossez et al. (2023). It consists of a combination of L1 and L2 distances between mel-spectrograms computed with different resolutions. For these experiments, we conducted a grid search for the weights of each reconstruction loss (details are in Section D.2) and report results for the best option in Table 4.
Table 4: Ablation study (VoxCeleb real data).
Loss
MOS ($\uparrow$)
UTMOS ($\uparrow$)
WV-MOS ($\uparrow$)
DNSMOS ($\uparrow$)
input $3.46\pm 0.07$ $2.77\pm 0.14$ $2.88\pm 0.16$ $2.72\pm 0.11$
2nd stage w/o reg. loss $3.71\pm 0.08$ $3.05\pm 0.09$ $2.70\pm 0.11$ $3.18\pm 0.06$
w/ L1Spec $4.15\pm 0.06$ $3.45\pm 0.08$ $3.41\pm 0.07$ $\textbf{3.28}\boldsymbol{\pm}\textbf{0.04}$
w/ RecLoss $4.07\pm 0.06$ $3.46\pm 0.07$ $3.43\pm 0.06$ $\textbf{3.28}\boldsymbol{\pm}\textbf{0.04}$
w/ LMOS $4.20\pm 0.05$ $3.48\pm 0.08$ $3.58\pm 0.06$ $3.26\pm 0.04$
+ WavLM enc. $4.21\pm 0.06$ $3.65\pm 0.08$ $3.75\pm 0.05$ $3.26\pm 0.04$
+ Scaling $4.30\pm 0.05$ $3.83\pm 0.07$ $\textbf{4.00}\boldsymbol{\pm}\textbf{0.06}$ $3.21\pm 0.05$
3rd st. + 3rd stage $4.59\pm 0.05$ $3.78\pm 0.07$ $\textbf{3.99}\boldsymbol{\pm}\textbf{0.06}$ $\textbf{3.29}\boldsymbol{\pm}\textbf{0.04}$
+ HF Loss $\textbf{4.63}\boldsymbol{\pm}\textbf{0.04}$ $\textbf{4.05}\pm\textbf{0.07}$ $\textbf{3.98}\boldsymbol{\pm}\textbf{0.06}$ $\textbf{3.31}\boldsymbol{\pm}\textbf{0.04}$
The other proposed improvements bring incremental gains in perceptual MOS quality. Firstly, we add the features of the WavLM encoder as an additional inputs to HiFi++. Next, we scale the architecture by increasing the number of channels within the model. After that, we concatenate the Upsample WaveUNet with the FINALLY-16 architecture and implement the 3rd stage of training on the DAPS-clean dataset (48 kHz). Finally, we use additional human feedback losses (HF Loss) during the 3rd stage to further improve perceptual quality.
Since Table 4 does not clearly demonstrate the advantages of using the WavLM (Chen et al., 2022b) encoder in terms of MOS, we have conducted additional ablation study on the more challenging UNIVERSE (Serrà et al., 2022) validation dataset. The results of these additional experiments, which are presented in Table 5, clearly demontarte the importance of WavLM encoder.
Table 5: Ablation of WavLM encoder on UNIVERSE validation data.
MOS ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) DNSMOS ($\uparrow$) PhER ($\downarrow$)
w/o WavLM $3.49\pm 0.08$ $3.33\pm 0.18$ $3.80\pm 0.15$ $\textbf{3.15}\boldsymbol{\pm}\textbf{0.09}$ $0.27\pm 0.04$
w/ WavLM $\textbf{3.75}\boldsymbol{\pm}\textbf{0.07}$ $\textbf{3.56}\boldsymbol{\pm}\textbf{0.20}$ $\textbf{3.99}\boldsymbol{\pm}\textbf{0.08}$ $\textbf{3.07}\boldsymbol{\pm}\textbf{0.16}$ $\textbf{0.21}\boldsymbol{\pm}\textbf{0.04}$
7 Conclusion
In conclusion, we theoretically demonstrate that LS-GAN training encourages the selection of the point of maximum density within the conditional clean speech distribution, aligning naturally with the objectives of speech enhancement. Our empirical investigation identifies WavLM as an effective backbone for perceptual loss, supporting adversarial training. By integrating WavLM-based perceptual loss into the MS-STFT adversarial training pipeline and enhancing the HiFi++ architecture with a WavLM encoder, we develop a novel speech enhancement model, FINALLY, which achieves state-of-the-art performance, producing clear and high-quality speech at 48 kHz.
Appendix A Additional results for perceptual losses
A.1 More on motivation
From the probabilistic point of view, minimization of the point-wise distance leads to an averaging effect. For example, optimization of the mean squared error between waveforms delivers the expectation of the waveform over the conditional distribution of clean speech given its degraded version $g_{\theta}(x)=\mathbb{E}_{y\sim p_{\text{clean}}(y|x)}[y]$. The key thing is that the expectation over the distribution is not guaranteed to lie in the regions of high density of this distribution (see Figure 4 and Figure 4).
Figure 3: Ground truth waveform and waveform resynthesized by HiFi GAN vocoder. While waveforms significantly differ, they correspond to the same sound, creating ambiguity in generation.
Figure 4: Regression in an entangled feature space might cause the expectation to lie outside the regions of high density, while regression in a disentangled space facilitates the expectation to lie within the regions of high probability density.
Mathematically speaking, let $\phi(x)$ be the mapping to this space, and assume there exists a reverse mapping $\phi^{-1}(z)$ such that $\phi^{-1}(\phi(x))=x$. We would like to note that in practice, for regression purposes, there is no need to explicitly know the reverse mapping $\phi^{-1}(z)$ as long as $\phi(x)$ is differentiable. The regression in the space produced by the mapping leads to averaging in this space; thus, after minimizing the MSE loss
$\mathbb{E}_{x,y\sim p_{\text{clean}}(y)p(x|y)}\|\phi(y)-\phi(g_{\theta}(x))\|^{2},$
one would obtain the solution
$g_{\theta}(x)=\phi^{-1}\left(\mathbb{E}_{y\sim p_{\text{clean}}(y|x)}[\phi(y)]\right).$
Therefore, a desirable property for the $\phi(x)$ mapping to be used as the regression space is that
[…]
i.e., we would like averaging in $\phi$-space to provide a representation corresponding to the most likely solution (the most probable reconstruction as discussed in the previous section). In practice, this property is difficult to verify. Moreover, finding such a mapping is likely to be a task which is not easier than the original problem. Therefore, based on this intuition, we propose some heuristic rules for assessing the structure of regression spaces produced by different mappings.
Figure 5: Comparison of training schemes on spectrograms. Going from the top to the bottom, the first spectrogram is obtained from the model trained only on WavLM (Chen et al., 2022b) features, the second one is produced by the model trained on both WavLM features and STFT L1 loss. The third spectrogram is obtained by training with both mentioned losses and adversarial loss. The last spectrogram is computed with ground truth audio.
A.2 Influence of losses on final audio
To illustrate the influence of LMOS loss, as well as the importance of adversarial training, we examine the impact of different training setup using vocoding and discuss the results using spectrograms in Figure 5. We start with training the vocoding model using only convolutional features from WavLM. This training turns out to be suboptimal, as the model produces noticeable artefacts, visible in the top spectrogram in Figure 5. Next, we complement the WavLM feature loss with the STFT L1 loss, which is a L1 distance between spectrograms of reference and predicted audios. This method yields better quality, however some artefacts still remain. Notice, that such training particularly struggles to reconstruct high frequencies and capture the harmonic content. Finally, we train the model using adversarial loss. For this experiment we use original HiFi GAN (Kong et al., 2020) setup. Adversarial training helps to alleviate artefacts and produces better quality, as compared to the regression training used in both former experiments. For these reasons, we rely both on our newly developed reconstruction loss and adversarial training for creating our speech enhancement model.
A.3 Comparison with prior work
We additionally compare the LMOS against other speech perceptual losses from the literature: (Hsieh et al., 2020; Close et al., 2023a; Defossez et al., 2020; Kong et al., 2020). The comparison is outlined in Table 6. The experiments are conducted on neural vocoding task using LJSpeech dataset (Ito & Johnson, 2017). The HiFi-GAN generator V1 was trained for 1M iterations with batch size 16 for each loss function.
Table 6: MOS on neural vocoding for different loss functions.
Method MOS ($\uparrow$)
PFPL (Hsieh et al., 2021) 2.40 $\pm$ 0.08
SSSR loss with HuBERT features (Close et al., 2023a) 2.58 $\pm$ 0.08
MS-STFT + L1 waveform (Defossez et al., 2020) 3.16 $\pm$ 0.08
LMOS (ours) 4.21 $\pm$ 0.07
adv. MPD-MSD (Kong et al., 2020) 4.65 $\pm$ 0.04
Ground Truth 4.66 $\pm$ 0.04
A.4 WavLM layers
To find suitable layer of WavLM (Chen et al., 2022b) model for our loss function, we take activations of each layer and measure Rand index and correlation with SNR. We seek to find such layer, that would produce high score for both metrics. Intuitively, this suggests that the layer’s activations create a space where audio with varying acoustics is effectively separated, and their hidden representations are responsive to noise.
Figure 6: Comparison of WavLM features using Rand index and negative correlation of SNR
As we see in Figure 6, the most suitable features come from the convolutional encoder or first transformer layer. For convenience, we compute loss using features from convolutional encoder (we call the these outputs WavLM-Conv features).
A.5 WavLM features visualization
To provide visual evidence for the successful choice of WavLM features, as well as to illustrate clustering rule and SNR rule, we show the first two principal components of WavLM-Conv features and depict corresponding clusters (Figure 7) and noisy features (Figure 8). Figure 7 illustrates that audio samples with the same lexical content form well-defined clusters, whereas those with different content are properly disentangled. Figure 8 shows that increasing SNR in audio moves it away for the centre of the cluster. These figures provide empirical evidence for the heuristics discussed in Section 3
Figure 7: Clustering rule visualized on Wavlm-Conv PCA features. The phrases corresponding to each cluster are visualized.
Figure 8: SNR rule visualized for Wavlm-Conv PCA features for a particular cluster.
Appendix B Comparison with MIIPHER
We compare the enhancement quality of LibriTTS test_other samples as released by the Miipher authors (Koizumi et al., 2023b). The results are outlined in Table 7. For the information on WER refer to Section D.1. computations.
Table 7: MOS Scores for comparison with Miipher (LibriTTS, test other).
Method MOS ($\uparrow$) UTMOS ($\uparrow$) DNSMOS ($\uparrow$) WV-MOS ($\uparrow$) WER ($\downarrow$)
Input 3.59 $\pm$ 0.07 3.41 $\pm$ 0.13 2.88 $\pm$ 0.12 3.69 $\pm$ 0.11 0.23 $\pm$ 0.04
Miipher 4.18 $\pm$ 0.06 3.95 $\pm$ 0.13 2.99 $\pm$ 0.12 4.15 $\pm$ 0.14 0.22 $\pm$ 0.05
Ours (16 kHz) 4.24 $\pm$ 0.05 4.18 $\pm$ 0.08 3.15 $\pm$ 0.09 4.28 $\pm$ 0.10 0.18 $\pm$ 0.04
Ours (48 kHz) 4.54 $\pm$ 0.05 4.18 $\pm$ 0.08 3.15 $\pm$ 0.09 4.28 $\pm$ 0.10 0.18 $\pm$ 0.04
Appendix C Implementation details
C.1 Reproducibility statement
Our model is based on HiFi++ architecture (Andreev et al., 2022), for which code is available in https://github.com/SamsungLabs/hifi_plusplus. We used several open-source datasets for training. More specifically, we used DAPS, LibriTTS-R and Deep Noise Suppression Challenge (DNS) (only noises). We rely on official implementations of MS-STFT discriminators, which can be found in https://github.com/facebookresearch/encodec. For LMOS loss and HiFi++ conditioining, we used WavLM large model from Hugging Face, which can be found in https://huggingface.co/microsoft/wavlm-large.
C.2 Augmentations
Before mixing with noise, we convolve the speech signal with a randomly chosen microphone impulse response from the Microphone Impulse Response project [^2] http://micirp.blogspot.com/ and apply other digital distortions. With a probability of 0.8, we also convolve the signal with a room impulse response randomly chosen from those provided in the DNS dataset. We additionally apply audio effects from the torchaudio library (Hwang et al., 2023), trying to simulate digital distortions. Parameter values are chosen randomly; only one codec is applied.
Table 8: Applied augmentations.
Augmentation Prob. Param. name Interval of values
acrusher 0.25 bits [1,9]
crystalizer 0.4 intensity [1,4]
flanger 0.15 depth [1,8]
vibrato 0.15 frequency [5,8]
codec ogg codec mp3 0.45 encoder vorbis, opus
bit rate [4000,16000]
C.3 Model architecture
In this section we detail parameters for our speech enhancement models. We use HiFi++ (Andreev et al., 2022) as backbone. This architecture consists of four parts: Spectral UNet, HiFi Upsampler, WaveUNet and SpectalMaskNet. In addition to that we also use WavLM-large feature encoder with transformer features, which takes waveform as an input and outputs last hidden state of the transformer, which is stacked with the SpectralUNet output and fed into the HiFi Upsampler. In the next paragraphs we thoroughly discuss the architecture of every part. Architecture summary is presented in Table 9
Residual blocks
To better understand Spectral UNet, WaveUNet and SpectralMaskNet we first discuss the key building block. Each residual block is composed of convolution, followed by LeakyReLU. We use weight norm (Salimans & Kingma, 2016) instead of BatchNorm, we found it to be more efficient in our preliminary experiments. We use kernel size 3 for 2D UNets and kernel size 5 for 1D UNet. Each residual block has additive skip connection. In our architectures we use a stacked composition of residual blocks. The number of residual blocks stacked together in one layer is called depth.
SpectralUNet architecture
SpectralUNet processes a mel-spectrogram input with dimensions [B, 80, T] and outputs a hidden state with dimensions [B, 512, T]. The input mel-spectrogram is combined with positional encoding. The architecture is based on a 2D UNet (Ronneberger et al., 2015) with five layers, each having respective channels of $[16,32,64,128,256]$ and a depth of 4. Additional convolutions are applied after these layers to transform 2D data from shape [B, 16, W, T] to [B, 1, W, T], and then to [B, 512, T]. This output is concatenated along the channel dimension with the final transformer hidden state of WavLM (Chen et al., 2022b), which has been interpolated using the nearest-neighbour method to match the output length of SpectralUNet in the time dimension. The concatenated result then passes through a Residual block with kernel size 3 and with a width of 1536 (1024 from the final transformer hidden state and 512 from SpectralUNet output),than passes through 1d Convolution with kernel size 1 and LeakyReLU to obtain size 512 in the channel dimension, and is subsequently fed into the HiFi Upsampler.
HiFi Upsampler architecture
For HiFi Upsampler we use HiFi generator architecture from (Kong et al., 2020). In our implementation we use 4 layer model with upsample rates $[8,8,2,2]$, kernel size $[16,16,4,4]$, hidden sizes $[512,256,128,64]$. For each layer we use residual blocks with kernels $[3,7,11]$ and dilations $[(1,3,5),(1,3,5),(1,3,5)]$.
WaveUNet architecture
We implement WaveUNet following (Stoller et al., 2018). We use 4-layer WaveUNet with channels $[128,128,256,512]$, each layer has depth 4. Our WaveUNet takes the concatenation of input waveform and HiFi Upsampler as input. In this way we provide a residual connection for Spectral UNet and HiFi Upsampler.
SpectralMaskNet architecture
The final Stage of our pipeline is SpectralMaskNet. It applies channel-wise Short Time Fourier Transform (STFT) to the input, decomposes it into phase and amplitude, and then processes amplitude with 2D UNet. Finally, it multiplies processed amplitude and phase and applies inverse STFT to obtain the final audio. For processing amplitude 2D UNet with channels $[64,128,256,512]$ as a backbone, each layer has depth 1. The output of SpectralMaskNet is the final output of our 16kHz model.
WaveUNet Upsampler
In order to enable the model to output samples in 48kHz we also add WaveUNet Upsampler, that upsamples the final audio. For that task we use WaveUNet with 5 layers and channels $[128,128,128,128,256]$, each layer has depth 3. The outputs of each layer are downsampled with the scale 4. After the output of the final upsample layer, we have an additional head that has 512 features and is used for directly upscaling the WaveUNet output into 48kHz.
Architecture summary
More details about the architecture can be found in (Andreev et al., 2022). We present the summary of parameters for Spectral UNet, WaveUNet and SpectralMaskNet in Table 9.
Table 9: Summary of parameters.
Submodel
Channel Sizes
Kernel Size
(constant for
each layer)
Layer Depth
(constant for
each layer)
Spectral UNet $[16,32,64,128,256]$ $3$ $4$
HiFi Upsampler $[512,256,128,64]$ $[3,7,11]$ $3$
WaveUNet $[128,128,256,512]$ $5$ $4$
Spectral MaskNet $[64,128,256,512]$ $3$ $1$
WaveUNet Upsampler $[128,128,128,128,256]$ + 512 $5$ $3$
Discriminators
We employ MS-STFT discriminators (Défossez et al., 2023) for adversarial training. We use 5 MS-STFT discriminators with n_ffts $[4096,2048,1024,512,256]$, hop_lengths $[1024,512,256,128,64]$, win_lengths $[4096,2048,1024,512,256]$ in 48kHz setting, for 16kHz one all above mentioned parameters are divided by 2. Our implementation closely follows (Défossez et al., 2023).
Training details
The main model is trained using 8 Nvidia P40 GPUs with effective batch size 32 and AdamW (Loshchilov & Hutter, 2017) as our main optimizer. For pretraining we use learning rate 0.0002, betas (0.8, 0.99) and learning rate exponential decay of 0.996 for each 200 iterations, the pretraining lasts 100,000 iterations. For the second stage we use learning rate 0.0002 with betas (0.8, 0.99) and learning rate decay 0.995, whereas for discriminators we use learning rate 0.0002 with betas (0.5, 0.999) and learning rate decay 0.995, the discriminators perform 2 optimization iterations for every 1 generator’s optimization iteration, the training lasts 30,000 generator’s iterations. We also use a linear warm-up for 2000 iterations for the generator. We use the same training parameters for the third stage during generator’s 40,000 iterations. The first and the second stage are trained in 16kHz using LibriTTS-R dataset with noises from DNS and augmentations, discussed in Footnote 2. Whereas the third stage is trained in 48kHz using DAPS dataset with noises from DNS as well. Datasets are discussed in Section C.1
Comparison of the datasets and resources used to train FINALLY and baseline models.
To better contextualize our results we provide a detailed comparison of the model sizes and training data of the baseline models.
Table 10: Comparison of resources and data used for training.
Model Training Data Scale (clean data) Model Size (parameters) RTF on V100 GPU
VoiceFixer 44 hours (VCTK) 112 M 0.02
DEMUCS 500 hours (DNS) 61 M 0.08
STORM 200 hours (WSJ0 and VCTK) 28 M 1.05
BBED 140 hours (WSJ0) 65 M 0.43
HIFI-GAN-2 5 hours (DAPS) 34 M 0.50
Universe 1500 hours (private data) 189 M 0.50
FINALLY (ours) 200 hours (LibriTTS-R and DAPS) 454 M (including 358 M of WavLM) 0.03
Although our model has more parameters compared to the baseline models, the majority of these parameters are dedicated to handling low-resolution features. For instance, the Transformer in WavLM processes representations of the waveform that are downsampled by a factor of 320, operating at 50 Hz. On the other hand, models such as HiFi-GAN-2 primarily work at the full waveform resolution, similar to WaveNet. This design enables our model to be more efficient in terms of computational resources, resulting in a significantly lower Real-Time Factor (RTF).
C.4 Limitations
The primary area for further developing in the proposed model lies in improving perceptual quality, particularly in instances when the speech is severely degraded. Although our model is capable of capturing the content of such signals, even in situations that may be impossible for humans, some artifacts remain. These artifacts could possibly be attributed to the high uncertainty of the initial signal, such as the speaker’s voice, for instance.
Another crucial aspect to consider is the streaming mode. While the proposed model is fast, it doesn’t suit low latency scenarios, which are widely applicable in telecommunications. The big obstacle to such improvement could be the WavLM (Chen et al., 2022b) model that utilizes context for identifying the content of speech. This model has the boundaries of the required context and, consequently, the minimum latency achievable.
Appendix D Evaluation details
D.1 Metric computation
To compare models in their ability to correctly reconstruct the speech we use Phoneme Error Rate (PhER) and Word Error Rate (WER). Given the reference audio and audio, produced by the speech enhancement model, metrics are computed in the following way. First, Automatic Speech Recognition (ASR) model is applied to both audios, yielding phoneme representations. Then, the editing distance is computed over predicted phonemes (or words for WER) for the reference and the predicted audio. The distance is normalized to the length of the reference phoneme representation (or the length of the sentence for WER) and averaged over the validation set. We rely on https://huggingface.co/ for the ASR models and metric computations. For PhER we use mrrubino/wav2vec2-large-xlsr-53-l2-arctic-phoneme model, for WER we use jonatasgrosman/wav2vec2-large-xlsr-53-english. These are Wav2Vec2-based (Baevski et al., 2020) models for ASR.
D.2 Ablation details
For the ablation study, we utilized a smaller version of our original model to expedite evaluations, with the reduced model containing approximately 22 million parameters. The dataset used for this ablation was provided by (Nagrani et al., 2017). The ablation process was carried out in several stages. Initially, we compared the performance of our LMOS loss with L1Spec loss (Kong et al., 2020), MS-STFT loss (Defossez et al., 2020), and RecLoss (Défossez et al., 2023) during the pretraining phase. Each loss function was used independently to train the model in a regressive manner for 500,000 steps with a batch size of 12, utilizing two Nvidia P40 GPUs. Our findings indicated that LMOS consistently outperformed the other losses. Detailed results are presented in Table 11. It’s important to note that MOS evaluations were not conducted during the first stage, as the enhanced audio typically displayed artifacts following this stage of training.
Table 11: Ablation for regression losses.
Metric UTMOS ($\uparrow$) WV-MOS ($\uparrow$) DNSMOS ($\uparrow$)
MS-STFT $2.54\pm 0.10$ $2.77\pm 0.10$ $3.04\pm 0.05$
RecLoss $2.53\pm 0.10$ $2.77\pm 0.10$ $3.04\pm 0.05$
L1Spec did not converge did not converge did not converge
LMOS $\mathbf{3.43\pm 0.09}$ $\mathbf{3.57\pm 0.05}$ $\mathbf{3.14\pm 0.04}$
Next, we ablated the use of mentioned losses in purely adversarial setup without pretraining. The same small model was trained from scratch using one of the discussed reconstruction losses with adversarial LS-GAN loss (Mao et al., 2017) and feature matching loss (Kong et al., 2020). We used MS-STFT discriminators (Défossez et al., 2023) for adversarial training. We fixed coefficients for the feature matching loss at 5 and for the LS-GAN loss at 15, and then grid searched the coefficient for each reconstruction loss. We experimented with the possible coefficient to be 1, 3, 10, 20, 50 and found that is all cases the best coefficient is 3. We found that LMOS loss outperforms other reconstruction losses in this setup as well.
In the first stage we experimented with the addition of SSL features to boost model quality. For these experiments we took our best model trained with LMOS loss. As our ablation indicates, transformer features of WavLM significantly increase the model quality.
Finally, we train our big model, using the ablated architecture. We report the results in ’Scaling’ row in Table 4. We also show how the introduction of the WaveUNet Upsampler and Human Feedback (HF) losses influence the final quality of our model.
In conclusion, we found that the best model is the one, trained with WavLM transformer features and LMOS reconstruction loss. Moreover, we found, that using HF losses also significantly increases the final quality.
Comparison with HiFi++.
Since our model is based on HiFi++ architecture (Andreev et al., 2022), use explicitly show how the improvements we introduced increase the quality of enhanced audio compared to HiFi++ baseline using VoxCelex dataset  (Nagrani et al., 2017). The results are presented in  Table 12
Table 12: Comparison with HiFi++ on VoxCeleb data.
Model UTMOS ($\uparrow$) WV-MOS ($\uparrow$) DNSMOS ($\uparrow$)
Input $2.72\pm 0.11$ $2.90\pm 0.16$ $2.72\pm 0.11$
HiFi++ $2.76\pm 0.13$ $2.68\pm 0.14$ $2.98\pm 0.07$
FINALLY (ours) $\mathbf{4.05\pm 0.07}$ $\mathbf{3.98\pm 0.06}$ $\mathbf{3.31\pm 0.04}$
Appendix E Subjective evaluation
Figure 9: The assessor’s interface.
Figure 10: The rules for the assessor.
We measure mean opinion score (MOS) of the model using a crowdsourcing adaptation of the standard absolute category rating procedure. Our MOS computing procedure is as follows.
  1. 1.
    Select a subset of 40 random samples from the test set (once per problem, i.e. for bandwidth extension or speech enhancement).
  2. 2.
    Select a set of models to be evaluated; inference their predictions on the selected subset.
  3. 3.
    Randomly mix the predictions and split them into the pages of size 20 almost uniformly. Almost uniformly means that on each page there are at least $\lfloor\frac{20}{\mathrm{num\_models}}\rfloor$ samples from each model.
  4. 4.
    Insert additional 4 trapping samples into random locations on each page: 2 samples from ground truth, and 2 samples of a noise without any speech.
  5. 5.
    Upload the pages to the crowdsourcing platform, set the number of assessors for each page to at least 30.
    Assessors are asked to work in headphones in a quiet environment; they must listen to the audio until the end before assess it.
  6. 6.
    Filter out the results where the ground truth samples are assessed with anything except 4 (good) and 5 (excellent), or the samples without voice are assessed with anything except 1 (bad).
  7. 7.
    Compute 95% confidence intervals via bootstrapping.
Since the models are distributed uniformly among the pages, assessor’s biases affect all models in the same way, so the relative order of the models remains. On the other hand, the assessor will have access to all variety of the models on one page and thus can scale his ratings better. The other side is that the models’ ratings depend on each other in this setting, because assessors tend to estimate the sample quality relatively to the average sample of the page. This means that when some models perform poorly in comparison, the higher scores are given to the better models. 4 trapping samples per page is also a reasonable choice, because one cannot just random guess the correct answers for these questions.
The instructions for the assessors and the screenshots of the evaluation interface are provided on Figure 9 and, Figure 10 respectively.
The total amount of money spent on the surveys while preparing the paper is somewhere between \$1000 and \$1500 (it is hard to estimate exactly because there were a lot of exploration tests during creating the model and writing the paper). According to the crowdsourcing system statistics, the average hourly wage varies between tasks from \$2.5 to \$4, which exceeds by a large margin the minimal wage in the countries where the test was conducted.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: -
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Appendix C.4
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer:[Yes]
  14. Justification: Please see the Appendix A
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Please see the Appendix C.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: We do not open the code for training of the model due to our organization policy, however, we use only publicly available data and most of the source codes are available as stated in Appendix C. Therefore, we expect that the work could be reproduced with moderate efforts.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Please see the Appendix C
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report the confidence intervals computed by bootstrapping.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: lease see the Appendix C
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: -
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: -
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: -
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: -
  60. Guidelines: Please the Section 5.
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: -
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: Please see the Appendix E
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: -
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.